Timberly Construction negotiates a lump-sum purchase of several assets from a company that is going out of business. The purchase is completed on January 1, 2017, at a total cash price of $830,000 for a building, land, land improvements, and four vehicles. The estimated market values of the assets are building, $494,000; land, $323,000; land improvements, $38,000; and four vehicles, $95,000. The company’s fiscal year ends on December 31. Required: 1-a. Prepare a table to allocate the lump-sum purchase price to the separate assets purchased. 1-b. Prepare the journal entry to record the purchase. 2. Compute the depreciation expense for year 2017 on the building using the straight-line method, assuming a 15-year life and a $32,000 salvage value. 3. Compute the depreciation expense for year 2017 on the land improvements assuming a five-year life and double-declining-balance depreciation.
Answer #1

1a) Table

Market values Percent of market value Lump sum price Allocated cost
Building 494000 52% 830000 431600
Land 323000 34% 830000 282200
Land improvement 38000 4% 830000 33200
Vehicle 95000 10% 830000 83000
Total 950000 100% 830000

1b) Journal entry

Date account and explanation Debit Credit
Building 431600
Land 282200
Land improvement 33200
Vehicle 83000
Cash 830000

2a) Depreciation expense for building = (431600-32000/15) = 26640

3) Depreciation expense for land improvement = 33200*40% = 13280